OpenStudy (anonymous):

Solve the equation for x. If necessary, round your answer to two decimal places. Do not include "x =" in your answer. ex = 27

OpenStudy (anonymous):

\[e^x = 27 \implies \ln 27 = x\] \[\ln 27 = 3.30\]

OpenStudy (anonymous):

oh thank you so much